# From Single Genes to Populations: Quantifying Broken Detailed Balance in Transcription

### v2

@@ II Entropy production of transcription in a two-state gene
The only data set to (i) split the data into cell cycle stages (removing some of the cell-to-cell variability), and (ii) exclude genes from the analysis that were not best modeled as a telegraph model is [19]. This leads to some discrepancies between the kinetic parameter distributions reported in [19] and [45, 46]. Here, we give more credence to results based on [19] due to the extra steps taken to ensure quality of the inferred parameters, but we still report results from the other datasets in the SI.
Figure 4: Investigating detailed balance breaking via $\dot{s}_{\mathrm{mes}}$ in real versus random combinations of experimental parameters found for G1 cell-cycle stage mouse fibroblasts in [19]. (a) Histogram over $\dot{s}_{\mathrm{mes}}$ for 1,436 real parameters sets (in gray), and 1,436 randomized parameters sets (in blue). Dashed lines show the respective mean values for each histogram. (b) Phase space plot of $\dot{s}_{\mathrm{mes}}$ versus $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ for real parameter sets. Each marker represents the position of a single gene in the phase space, and the $r$ value in the plot title is Pearson’s correlation coefficient between the logarithm of the $x$ and $y$ variables. (c) Phase space plot of $\dot{s}_{\mathrm{mes}}$ versus $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ for randomized parameter sets. (d) Phase space plot of $\dot{s}_{\mathrm{mes}}$ versus $CV^{2}$ for real parameter sets. (e) Phase space plot of $\dot{s}_{\mathrm{mes}}$ versus $CV^{2}$ for randomized parameter sets.
@@
All of these datasets report values of $\{\rho_{\mathrm{on}},\sigma_{\mathrm{off}},\sigma_{\mathrm{on}}\}$. To obtain values for the mRNA degradation rate, $d$, we crossed referenced gene names with the mouse mRNA degradation rate dataset from [71]. Note that these degradation rates were measured in mouse embryonic stem cells, and here we assume that they are a good approximation of the degradation rates of mRNA in mouse fibroblast cells. Finally, since values of $\rho_{\mathrm{off}}$ have not been inferred (due to parameter indistinguishability), we set $\rho_{\mathrm{off}}=10^{-3}s^{-1}$. Note that this value of $\rho_{\mathrm{off}}$ is quite arbitrary as long as it is chosen such that $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$ for two reasons. First, the logarithmic dependence of $\dot{s}_{\mathrm{mes}}$ on $\rho_{\mathrm{off}}$ implies that making $\rho_{\mathrm{off}}$ orders of magnitude smaller only changes $\dot{s}_{\mathrm{mes}}$ by a multiplicative factor of $\mathcal{O}(1)$. In our calculation of the lower bound on the EPR of mRNA transcription below, this means that $\dot{s}_{\mathrm{mes}}$ is relatively insensitive the value of $\rho_{\mathrm{off}}$, so long as it is chosen such that $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$. Second, in our subsequent analyses we either: (i) compare the patterns seen in the real parameter sets to randomly drawn parameter sets, but for the same value of $\rho_{\mathrm{off}}$ (in Fig. 4); or (2) compare values of $\dot{s}_{\mathrm{mes}}$ across parameter sets where the interest is in the differences in $\dot{s}_{\mathrm{mes}}$ across parameter space.
In Fig. 4, we compare $\dot{s}_{\mathrm{mes}}$ calculated from real parameter sets for mouse fibroblast cells in the G1 cell-cycle stage [19] to $\dot{s}_{\mathrm{mes}}$, calculated from random parameter sets. The random parameter sets were constructed from the real dataset such that each parameter set $\{\rho_{\mathrm{on}},\sigma_{\mathrm{off}},\sigma_{\mathrm{on}},d\}$ has kinetic parameters that are drawn uniformly (with replacement) from the real values of those parameters.
@@
We find several key features, that are broadly exhibited across all seven datasets in Figs. S2-S7. First, as shown in Fig. 4(a), the real parameter sets have EPRs that cluster much more tightly than the randomized parameter sets, meaning that real gene expression is characterized by a more tightly regulated EPR. Second, there is a systematically lower value mean of $\dot{s}_{\mathrm{mes}}$ for the real parameter sets (shown by the black dashed line) than the randomized parameters (shown by the blue dashed line). This indicates that real transcription is somehow benefited by having a reduced $\dot{s}_{\mathrm{mes}}$ than would randomly occur. Third, we looked at $\dot{s}_{\mathrm{mes}}$ against $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$. Comparing Fig. 4(b) and Fig. 4(c), we find that the real phase space in Fig. 4(b) consists of two distinct clusters with an overall small correlation coefficient (although the right-hand cluster has a clearly has a negative correlation), while the random parameter sets exhibit a stronger negative correlation without the clear demarcation of two distinct clusters. For real parameter sets, these results highlight that there are two distinct types transcriptional behavior, and that in the burstier of these groups that increased burstiness corresponds to decreased $\dot{s}_{\mathrm{mes}}$. Finally, we use the formula for the coefficient of variation squared in Eq. (S4) to investigate the relationship between $\dot{s}_{\mathrm{mes}}$ and mRNA noise. Comparing Fig. 4(d) and Fig. 4(e), we show that real parameter sets correspond to an EPR that is correlates negatively with mRNA expression noise. This correlation is not observed for randomized parameter sets. In general, more bursty behaviors correspond to reduced values of $\dot{s}_{\mathrm{mes}}$.
@@
Figure 5: Assessing the dependence of $\dot{s}_{\mathrm{mes}}$ against kinetic parameters across 1,436 genes, where $\langle\dot{s}_{\mathrm{mes}}\rangle$ is the mean EPR for binned values of the kinetic parameters. In all plots, pink lines correspond to the mode of the frequency distribution over a given kinetic parameter. (a) Binning genes across $\log_{10}(B)$ shows an increasing dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ on $\log_{10}(B)$. The modal value of $\log_{10}(B)$ occurs well before the rapid increase of $\langle\dot{s}_{\mathrm{mes}}\rangle$. (b) Binning genes across $\log_{10}(\sigma_{\mathrm{on}})$, with similar conclusions as for plot (a). (c) Binning genes across $\log_{10}(\sigma_{\mathrm{off}})$ shows a surprising feature—that the modal values of $\log_{10}(\sigma_{\mathrm{off}})$ occur before and after the peak in $\langle\dot{s}_{\mathrm{mes}}\rangle$. (d) For $\log_{10}(\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}})$ similar conclusions are reached as for panel (c).
@@
To investigate the average dependence of $\dot{s}_{\mathrm{mes}}$ for different values of kinetic parameters, we explore at the dependence of $\dot{s}_{\mathrm{mes}}$ binned across kinetic parameters. To perform this binning procedure, we take one of the kinetic parameters, say $B$, split it into $M$ equally spaced bins (in log-space) and then average the corresponding values of $\dot{s}_{\mathrm{mes}}$ in each bin to give,
@@
In Fig. 5, we show the binning-based analysis on G1 cell-cycle stage mouse fibroblasts. Each panel shows (above) the binned $\langle\dot{s}_{\mathrm{mes}}\rangle$ values across the measured range of kinetic parameters, and (below) the frequency distribution of parameter values. This allows for a side-by-side comparison to observe whether there is a relationship between $\langle\dot{s}_{\mathrm{mes}}\rangle$ and the distribution of parameter values.
@@
There are several key findings for across the four panels shown in Fig. 5—and across all seven datasets (see Figs. S9-S14)—centered around a common theme: that peaks in the frequency distributions of kinetic parameters describe areas of low $\langle\dot{s}_{\mathrm{mes}}\rangle$. This hints that having a low value of $\langle\dot{s}_{\mathrm{mes}}\rangle$ may be evolutionarily beneficial for transcription. This is most strikingly seen in Figs. 5(c) and (d), where the peaks of the frequency distribution are found on either side of a peak in $\langle\dot{s}_{\mathrm{mes}}\rangle$. This same feature is also observed for G2M cell-cycle stage mouse fibroblasts in Fig. S9(c). To assess whether the choice of averaging of $\dot{s}_{\mathrm{mes}}$ affects this observed phenomenology, we explored using the median, minimum and maximum of $\dot{s}_{\mathrm{mes}}$ in a given bin. In Fig. S8, we show that using these other metrics recapitulates the same phenomenon, and the peak in a given metric does not coincide with a peak in the frequency distribution. For datasets from [45] and [46] (see Figs. S10-S14) the range of values of $\langle\dot{s}_{\mathrm{mes}}\rangle$ found across each kinetic parameter was found to be far less that the dataset in [19], meaning that the phenomenon of high parameter abundance implying low $\langle\dot{s}_{\mathrm{mes}}\rangle$ is not quite as pronounced.
@@
These surprising findings point towards large values of $\dot{s}_{\mathrm{mes}}$ being disfavored for a genes transcriptional kinetics. One could think of this as a mesoscopic analog of energy expenditure minimization [72] or the principle of least dissipation [73], both of which at a microscopic level state that to be energy efficient is more beneficial. So far, $\dot{s}_{\mathrm{mes}}$ has been interpreted as a measure of irreversibility, and not as entropy production in the thermodynamic sense. Below, we not assess whether these results have any thermodynamic relevance.
@@
There is a large literature that discusses how coarse-grained systems provide lower bounds on the EPR of the corresponding microscopic system [30, 31, 32, 2, 33, 34, 35]. Using these references, in combination with previous work on housekeeping free energy usage [64], it follows that
@@
where $\overline{f_{\mathrm{dis}}}$ is the average rate of free energy dissipation per cell necessary to keep the system out of equilibrium, and $T$ is the temperature of the cell. Here, if we assume that the two-state gene model is a coarse-grained model of all the microscopic processes making up the transcription of a single gene, then Eqs. (10) and (11) should provide a lower bound on the free energy rate necessary to keep a gene transcribing with parameters $\{\rho_{\mathrm{on}},\rho_{\mathrm{off}},\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},d\}$. Based on England’s calculations in [2], given that a coarse-grained birth-death process seems provides a relatively tight lower bound on the thermodynamics of a proliferating population of cells, we may be cautiously optimistic that the two-stage model of gene expression can provide a lower energy bound for the transcription of a single gene.
To assess the tightness of $T\dot{s}_{\mathrm{mes}}$ as a lower bound on $\overline{f_{\mathrm{dis}}}$, we can make a crude order of magnitude estimate on the energy needed for transcription based on measured values. Let us say that the energy needed for transcription is in the turnover of producing mRNA transcripts from pre-existing nucleotides. The energy cost of this is 2 phosphorylation events per chain elongation step (i.e., per nucleotide) [74], and the energy of a single phosphorylation event is around $10kT$ [75, Sec. What is the energy in transfer of a phosphate group?]. Therefore, for a gene of length $L$, with kinetic parameters $\{\rho_{\mathrm{on}},\rho_{\mathrm{off}},\sigma_{\mathrm{on}},\sigma_{\mathrm{off}}\}$ (where $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$) the energy usage per second is approximately
@@
For a gene of length $1k\cdot bp$, and mean values of $\sigma_{\mathrm{on}}=0.47\,s^{-1},\sigma_{\mathrm{off}}=77.25\,s^{-1},\rho_{\mathrm{on}}=62.85\,s^{-1}$ (from G1 cell-cycle stage mouse fibroblast cells) this results in $\tilde{f}\approx 3800kT$. Given that the largest values of $\dot{s}_{\mathrm{mes}}$ for the same dataset are of order $\mathcal{O}(10^{2})$ (see Fig. 4), this shows that one cannot interpret $\dot{s}_{\mathrm{mes}}$ as having thermodynamic relevance. Recent work in [37] has shown that the results of England in [2] should be exercised with caution. The results of this section are in line with these findings.
III Entropy production of transcription in a population of cells
In the previous section, we focused on the EPR of transcription of a single gene assuming a canonical two-state gene. Scaling-up this calculation up for a population of identical cells is straightforward—the EPR is simply multiplied by the number of cells $N$ (if the cells are assumed to be non-interacting). However, populations of cells vary in many features, such as cell size, pH and morphology. Therefore, inferring a single set of kinetic parameters for a population with variable kinetics can lead to significant inferential error [24, 76] (see Fig. 1(b)). By extension, the subtleties of extrinsic noise may lead to properties of reversibility of an ensemble that are non-trivially related to those of a single cell.
Following [24], we assume that the kinetic rates of the two-state model are gamma distributed—since the mean and variance of the distribution can be tuned independently and the gamma distribution is defined for positive real numbers (kinetic rates are, by definition, positive). Then, each cell in the population has kinetic parameters which are drawn from these distributions. We investigate the effect of having this extrinsic noise parameter-by-parameter, for two reasons: (i) it allows for the isolation of the effects of noise on that parameter on the EPR and (ii) it allows for analytic results for $\dot{s}_{\mathrm{mes}}$ across a variable population. $\theta$ represents one of $\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},d$ and $\rho_{\mathrm{off}}$ in a single cell. The gamma distribution has the probability density function
@@ III Entropy production of transcription in a population of cells
where $\alpha$ represents the shape parameter and $\beta$ the inverse scale parameter, in which the mean is $\overline{\theta}=\alpha/\beta$ and the variance as $\mathrm{Var}(\theta)=\alpha/\beta^{2}$. We use $\overline{x}$ to denote the average of $x$ over the population. For comparison with the single cell case, $\dot{s}_{\mathrm{mes}}[\theta]$ is the average mesoscopic EPR per cell, an intrinsic quantity, given by
$\dot{s}_{\mathrm{mes}}[\theta;\alpha,\beta]=\int_{0}^{\infty}\frac{f(\theta;\alpha,\beta)(\rho_{\mathrm{on}}-\rho_{\mathrm{off}})\sigma_{\mathrm{on}}\sigma_{\mathrm{off}}}{(\sigma_{\mathrm{on}}+\sigma_{\mathrm{off}})(d+\sigma_{\mathrm{on}}+\sigma_{\mathrm{off}})}\ln\left(\frac{\rho_{\mathrm{on}}}{\rho_{\mathrm{off}}}\right)\mathrm{d}\theta,$ (15)
where the functional parameter $\theta\in\{\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},d,\rho_{\mathrm{off}}\}$ denotes the parameter that is varying over the population. We also define

### v3

@@ Introduction
Only the data sets from [sukys2025cell]: (i) split the data into cell-cycle stages (removing some of the cell-to-cell variability); and (ii) excluded genes from the analysis that were not best modeled as a telegraph model. This leads to some discrepancies between the kinetic parameter distributions reported in [sukys2025cell] and [larsson2019genomic, ramskold2024single]. Notably, in the data of [larsson2019genomic, ramskold2024single] there is a hard inferential lower bound imposed on the burst size at $B=1$, which may not be a good boundary choice for many genes. Here, we give more credence to results based on [sukys2025cell] due to the extra steps taken to ensure quality of the inferred parameters, but we still report results from the other datasets in the SI.
@@
All of these datasets report values of $\{\rho_{\mathrm{on}},\sigma_{\mathrm{off}},\sigma_{\mathrm{on}}\}$. To obtain values for the mRNA degradation rate, $d$, we cross-referenced gene names with the mouse mRNA degradation rate dataset from [sharova2009database]. Note that these degradation rates were measured in mouse embryonic stem cells, and here we assume that they are also a good approximation of the degradation rates of mRNA in mouse fibroblast cells. Finally, since values of $\rho_{\mathrm{off}}$ have not been inferred (due to parameter indistinguishability), we set $\rho_{\mathrm{off}}=10^{-3}s^{-1}$. Note that this value of $\rho_{\mathrm{off}}$ is quite arbitrary as long as it is chosen such that $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$ for two reasons. First, the logarithmic dependence of $\dot{s}_{\mathrm{mes}}$ on $\rho_{\mathrm{off}}$ implies that making $\rho_{\mathrm{off}}$ orders of magnitude smaller only changes $\dot{s}_{\mathrm{mes}}$ by a multiplicative factor of order $\mathcal{O}(1)$. In our calculation of the lower bound on the EPR of mRNA transcription below, this means that $\dot{s}_{\mathrm{mes}}$ is relatively insensitive the value of $\rho_{\mathrm{off}}$, so long as it is chosen such that $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$. Second, in our subsequent analyses we either: (i) compare the patterns seen in the real parameter sets to randomly drawn parameter sets, but for the same value of $\rho_{\mathrm{off}}$ (in Fig. 4); or (ii) compare values of $\dot{s}_{\mathrm{mes}}$ across parameter sets where the interest is in the differences in $\dot{s}_{\mathrm{mes}}$ across parameter space.
Statistical Null Models
For comparison to the real kinetic parameters sets introduced above, we introduce two different null models that serve as “neutral” comparisons when making claims about how physically realized parameter sets may impact the EPR.
Null model 1: To assess the degree to which specific combinations of real parameters affect the EPR we introduce a null model in which parameters are independently randomly sampled from the experimental values (with replacement) for each gene in the dataset. This ensemble is then comprised of the same number of “genes” but where all interdependencies between kinetic parameters have been removed. Therefore, comparisons between the real data and null model 1 test whether specific combinations of parameters have an influence on the magnitude of $\dot{s}_{\mathrm{mes}}$, isolating the influence of the combinations of physically realized parameters on the EPR.
Null model 2: To assess the degree to which specific kinetic parameters affect the EPR we introduce a null model in which the kinetic parameters themselves are randomly sampled from neutral probability distributions (i.e., not from the data). Because each kinetic parameter varies over several orders of magnitude, we sampled each parameter from a log-uniform distribution where the interval of the distribution is defined by the minimum and maximum values of each kinetic parameter in the data. This ensemble is also comprised of the same number of “genes” but where the kinetic parameters are randomly sampled. Therefore, comparisons between the real data and null model 2 test whether the specific values of kinetic parameters have significant influence over $\dot{s}_{\mathrm{mes}}$, or whether Eq. (11) and similar order of magnitude values of the kinetic parameters are enough to explain the magnitude of $\dot{s}_{\mathrm{mes}}$. That is, null model 2 isolates the impact of the distribution of kinetic parameters on the EPR.
Figure 4: Investigating detailed balance breaking via $\dot{s}_{\mathrm{mes}}$ between real parameters sets and null models 1 and 2 for G1 cell-cycle stage mouse fibroblasts in [sukys2025cell]. (a) Histogram over $\dot{s}_{\mathrm{mes}}$ for 1,436 real parameters sets (in gray), 1,436 parameter sets under null model 1 (in blue), and 1,436 parameter sets under null model 2 (in pink). Dashed lines show the respective mean values for each histogram. (b) Plot of $\dot{s}_{\mathrm{mes}}$ versus $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ for real parameter sets. Each marker represents the position of a single gene in the phase space, and the $r$ value in the plot title is Pearson’s correlation coefficient between the logarithm of the $x$ and $y$ variables. (c)-(d) Plots of $\dot{s}_{\mathrm{mes}}$ versus $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ under null models 1 and 2 respectively. (e) Plot of $\dot{s}_{\mathrm{mes}}$ versus $CV^{2}$ for real parameter sets. (f)-(g) Plots of $\dot{s}_{\mathrm{mes}}$ versus $CV^{2}$ under null models 1 and 2 respectively.
Relationship between $\dot{s}_{\mathrm{mes}}$ and real parameters
@@
In Fig. 4, we compare $\dot{s}_{\mathrm{mes}}$ calculated from real parameter sets for mouse fibroblast cells in the G1 cell-cycle stage [sukys2025cell] to $\dot{s}_{\mathrm{mes}}$ evaluated under the null models. We find several key features, that are broadly exhibited across all seven datasets in Figs. S2-S7. First, as shown in Fig. 4(a), the real parameter sets have EPRs that cluster much more tightly than the null models, meaning that real gene expression is characterized by a more tightly regulated EPR. Second, there is a systematically lower value mean of $\dot{s}_{\mathrm{mes}}$ for the real parameter sets than the null models. This indicates: (i) that transcription has kinetics that have lower $\dot{s}_{\mathrm{mes}}$ than occurs through randomly assigned but real kinetic values (null model 1) and (ii) that the real values of kinetic parameters reduce the EPR compared to kinetic parameters that are fully randomized (null model 2). Third, we looked at $\dot{s}_{\mathrm{mes}}$ against $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$. Comparing Figs. 4(b)-(d), we find that the real parameter sets lead to two distinct clusters with an overall small correlation coefficient (although the right-hand cluster has a clearly has a negative correlation), while the null models exhibit a stronger negative correlation without the clear demarcation of two distinct clusters. For real parameter sets, these results highlight that there are two distinct types transcriptional behavior, and in the burstier of these groups that increased burstiness corresponds to decreased $\dot{s}_{\mathrm{mes}}$. Finally, we use the formula for the coefficient of variation squared in Eq. (S4) to investigate the relationship between $\dot{s}_{\mathrm{mes}}$ and mRNA noise. Comparing Figs. 4(e)-(g), we show that real parameter sets correspond to an EPR that is correlates negatively with mRNA expression noise. This correlation is weaker, albeit still present, for randomized parameter sets. In general, burstier behaviors correspond to reduced values of $\dot{s}_{\mathrm{mes}}$.
@@
Figure 5: Assessing the dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ against real kinetic parameter combinations across 1,436 genes for G1 cell-cycle stage mouse fibroblasts in [sukys2025cell]. Here, $\langle\dot{s}_{\mathrm{mes}}\rangle$ is the mean EPR for binned values of real kinetic parameters sets. In all plots, pink lines correspond to the mode of the frequency distribution over a given kinetic parameter. (a) Binning genes across $\log_{10}(B)$ shows an increasing dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ on $\log_{10}(B)$. The modal value of $\log_{10}(B)$ occurs well before the rapid increase of $\langle\dot{s}_{\mathrm{mes}}\rangle$. (b) Binning genes across $\log_{10}(\sigma_{\mathrm{on}})$, with similar conclusions as for plot (a). (c) Binning genes across $\log_{10}(\sigma_{\mathrm{off}})$ shows a surprising feature—that the modal values of $\log_{10}(\sigma_{\mathrm{off}})$ occur before and after the peak in $\langle\dot{s}_{\mathrm{mes}}\rangle$. (d) For $\log_{10}(\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}})$ similar conclusions are reached as for panel (c).
@@
To investigate the average dependence of $\dot{s}_{\mathrm{mes}}$ for different values of kinetic parameters, we explore the dependence of $\dot{s}_{\mathrm{mes}}$ binned across kinetic parameters. To perform this binning procedure, we take one of the kinetic parameters, say $B$, split it into $M$ equally spaced bins (in log-space) and then average the corresponding values of $\dot{s}_{\mathrm{mes}}$ in each bin to give,
@@
In Fig. 5, we show the binning-based analysis on G1 cell-cycle stage mouse fibroblasts. Each panel shows (above) the binned $\langle\dot{s}_{\mathrm{mes}}\rangle$ values across the measured range of kinetic parameters, and (below) the frequency distribution of parameter values. This allows for a side-by-side comparison to observe whether there is a relationship between $\langle\dot{s}_{\mathrm{mes}}\rangle$ and the distribution of parameter values. In Figs. 6 and 7 we replicate the same figure but for null models 1 and 2 respectively.
Figure 6: Assessing the dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ against realizations of null model 1 for G1 cell-cycle stage mouse fibroblasts in [sukys2025cell]. (a) Binning genes across $\log_{10}(B)$ shows little difference to Fig. S1(a). (b) Binning across $\log_{10}(\sigma_{\mathrm{on}})$ now results in the modal value of $\sigma_{\mathrm{on}}$ being located at higher $\langle\dot{s}_{\mathrm{mes}}\rangle$ than for the real kinetic parameter combinations. (c) Binning across $\log_{10}(\sigma_{\mathrm{off}})$ shows one mode coinciding with a peak in the $\langle\dot{s}_{\mathrm{mes}}\rangle$. (d) For $\log_{10}(\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}})$ similar conclusions are reached as for panel (c).
Figure 7: Assessing the dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ against realizations of null model 2 for G1 cell-cycle stage mouse fibroblasts in [sukys2025cell]. Across (a)-(d) we see little relationship between $\langle\dot{s}_{\mathrm{mes}}\rangle$ and kinetic parameters.
@@
There are several key findings for across Figs. 5-7 centered around a common theme: that peaks in the frequency distributions of kinetic parameters describe areas of low $\langle\dot{s}_{\mathrm{mes}}\rangle$. This hints that low values of $\langle\dot{s}_{\mathrm{mes}}\rangle$ are characteristic of real transcription. This is most strikingly seen in Figs. 5(c) and (d), where the peaks of the frequency distribution are found on either side of a peak in $\langle\dot{s}_{\mathrm{mes}}\rangle$. This same feature is also observed for G2M cell-cycle stage mouse fibroblasts in Fig. S9(c). To assess whether the choice of averaging of $\dot{s}_{\mathrm{mes}}$ affects this observed phenomenology, we explored using the median, minimum and maximum of $\dot{s}_{\mathrm{mes}}$ in a given bin. In Fig. S8, we show that using these other metrics recapitulates the same phenomenon, and the peak in a given metric does not coincide with a peak in the frequency distribution.
For the null models, areas of low $\langle\dot{s}_{\mathrm{mes}}\rangle$ are not necessarily indicative of peaks in kinetic parameter values. For null model 1, we see in Fig. 6(b) that the modal value of $\sigma_{\mathrm{on}}$ no longer precedes the increase in $\langle\dot{s}_{\mathrm{mes}}\rangle$. Figs. 6(c)-(d) show that although the more pronounced of the two modes is located in an area of relatively low $\langle\dot{s}_{\mathrm{mes}}\rangle$ the wider left-hand mode is located close to the peak in $\langle\dot{s}_{\mathrm{mes}}\rangle$—with a value of $\langle\dot{s}_{\mathrm{mes}}\rangle$ approximately 100 times higher than for the real parameter sets. This indicates that in the burstier regime of $\sigma_{\mathrm{off}}\gg\sigma_{\mathrm{on}}$ that $\langle\dot{s}_{\mathrm{mes}}\rangle$ will be low regardless of the values of $B$ and $d$, whereas genes for which $\sigma_{\mathrm{off}}\approx\sigma_{\mathrm{on}}$ obtain low $\langle\dot{s}_{\mathrm{mes}}\rangle$ by varying the entire combination of parameter values, including $B$ and $d$. Note that the values of $\langle\dot{s}_{\mathrm{mes}}\rangle$ are in general higher for null model 1 than the real kinetic parameter sets as shown in Fig. 4(a). For null model 2, we see that fully randomized parameters remove the dependence of $\langle\dot{s}_{\mathrm{mes}}\rangle$ on the kinetic parameters. We observe the same properties of the null models for G2M cell-cycle stage mouse fibroblast genes in Figs. S15 and S16.
@@
These findings point towards large values of $\dot{s}_{\mathrm{mes}}$ being disfavored for a genes transcriptional kinetics. One could think of this as a mesoscopic analog of energy expenditure minimization [govern2014optimal] or the principle of least dissipation [han2008least], both of which at a microscopic level state that to be energy efficient is more beneficial. For mesoscopic systems, such principles have been hinted to in recent work [ray2025large]. So far, $\dot{s}_{\mathrm{mes}}$ has been interpreted as a measure of irreversibility, and not as an entropy production rate in the thermodynamic sense. In the next section we assess whether these results have any thermodynamic relevance.
For datasets from [larsson2019genomic] and [ramskold2024single] (see Figs. S10-S14) the range of values of $\langle\dot{s}_{\mathrm{mes}}\rangle$ found across each kinetic parameter was found to be far less that the dataset in [sukys2025cell], meaning that the phenomenon of high parameter abundance implying low $\langle\dot{s}_{\mathrm{mes}}\rangle$ is not an obvious conclusion. We suspect that the imposition of the inferential lower bound $B=1$ has some influence on the differing conclusions between the data from [sukys2025cell] and those from [larsson2019genomic, ramskold2024single].
@@
There is a large literature that discusses how coarse-grained systems provide lower bounds on the EPR of the corresponding microscopic system [egolf2000equilibrium, blythe2008reversibility, gomez2008lower, england2013statistical, bo2017multiple, jia2016model, skinner2021estimating]. Using these references, in combination with previous work on housekeeping free energy usage [ge2013dissipation], it follows that
@@
where $\overline{f_{\mathrm{dis}}}$ is the average rate of free energy dissipation per cell necessary to keep the system out of equilibrium, and $T$ is the temperature of the cell. Here, if we assume that the two-state gene model is a coarse-grained model of all the microscopic processes making up the transcription of a single gene, then Eqs. (10) and (11) should provide a lower bound on the free energy rate necessary to keep a gene transcribing with parameters $\{\rho_{\mathrm{on}},\rho_{\mathrm{off}},\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},d\}$. Based on England’s calculations in [england2013statistical], given that a coarse-grained birth-death process seems provides a relatively tight lower bound on the thermodynamics of a proliferating population of cells, we may be cautiously optimistic that the two-stage model of gene expression can provide a lower energy bound for the transcription of a single gene.
To assess the tightness of $T\dot{s}_{\mathrm{mes}}$ as a lower bound on $\overline{f_{\mathrm{dis}}}$, we can make an order of magnitude estimate on the energy needed for transcription based on measured values. Let us say that the energy needed for transcription is in the turnover of producing mRNA transcripts from pre-existing nucleotides. The energy cost of this is 2 phosphorylation events per chain elongation step (i.e., per nucleotide) [lynch2015bioenergetic], and the energy of a single phosphorylation event is around $10kT$ [milo2015cell, Sec. What is the energy in transfer of a phosphate group?]. Therefore, for a gene of length $L$, with kinetic parameters $\{\rho_{\mathrm{on}},\rho_{\mathrm{off}},\sigma_{\mathrm{on}},\sigma_{\mathrm{off}}\}$ (where $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$) the energy usage per second is approximately
@@
For a gene of length $1k\cdot bp$, and mean values of $\sigma_{\mathrm{on}}=0.47\,s^{-1},\sigma_{\mathrm{off}}=77.25\,s^{-1},\rho_{\mathrm{on}}=62.85\,s^{-1}$ (from G1 cell-cycle stage mouse fibroblast cells) this results in $\tilde{f}\approx 3800kT$. Given that the largest values of $\dot{s}_{\mathrm{mes}}$ for the same dataset are of order $\mathcal{O}(10^{3})$ with a mean of order $\mathcal{O}(10^{1})$ (see Fig. 4), this shows that one cannot interpret $\dot{s}_{\mathrm{mes}}$ as having thermodynamic relevance. Recent work in [kolchinsky2024thermodynamic] has shown that the results of England in [england2013statistical] should be exercised with caution. The results of this section are in line with these findings.
Entropy production of transcription in a population of cells
In the previous section, we focused on the EPR of transcription of a single gene assuming a canonical two-state gene. Scaling-up this calculation up for a population of identical cells is straightforward—the EPR is simply multiplied by the number of cells $N$ (if the cells are assumed to be non-interacting). However, populations of cells vary in many features, such as cell size, pH and morphology. Therefore, inferring a single set of kinetic parameters for a population with variable kinetics can lead to significant inferential error [lim2015quantitative, grima2023quantifying, gorin2024new] (see Fig. 1(b)). By extension, the subtleties of extrinsic noise may lead to properties of reversibility of an ensemble that are non-trivially related to those of a single cell.
Following [ham2020extrinsic, grima2023quantifying], we assume that the kinetic rates of the two-state model are gamma distributed—since the mean and variance of the distribution can be tuned independently and the gamma distribution is defined for positive real numbers (kinetic rates are, by definition, positive). Then, each cell in the population has kinetic parameters which are drawn from these distributions. We investigate the effect of having this extrinsic noise parameter-by-parameter, for two reasons: (i) it allows for the isolation of the effects of noise on that parameter on the EPR and (ii) it allows for analytic results for $\dot{s}_{\mathrm{mes}}$ across a variable population. Below, $\theta$ represents one of $\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},d$ and $\rho_{\mathrm{off}}$ in a single cell. The gamma distribution has the probability density function
@@
where $\alpha$ represents the shape parameter and $\beta$ the inverse scale parameter, in which the mean is $\overline{\theta}=\alpha/\beta$ and the variance as $\mathrm{Var}(\theta)=\alpha/\beta^{2}$. We use $\overline{x}$ to denote the average of $x$ over a population of cells. For comparison with the single cell case, $\dot{s}_{\mathrm{mes}}[\theta]$ is the average mesoscopic EPR per cell, an intrinsic quantity, given by
$\dot{s}_{\mathrm{mes}}[\theta;\alpha,\beta]=\int_{0}^{\infty}\dot{s}_{\mathrm{mes}}f(\theta;\alpha,\beta)\mathrm{d}\theta,$ (15)
where $\dot{s}_{\mathrm{mes}}$ in the integrand is taken from Eq. (10), and the functional parameter $\theta\in\{\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},d,\rho_{\mathrm{off}}\}$ denotes the parameter that is varying over the population. We also define